# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

### v4

@@ 4 Experiment
Losses design. We conduct ablation studies on the MSR-VTT 1k-A test set to assess the effectiveness of the proposed increment $\Delta$ and its associated regularizers. As shown in the right of Table 3, directly injecting $\Delta$ into the InfoNCE flow improves performance from 46.6 to 47.4, validating the benefit of gradient tension release via pairwise adjustment. Introducing the information bottleneck (IB) loss further boosts performance to 48.3, highlighting its role in guiding $\Delta$ toward semantically meaningful corrections. However, adding the norm constraint or directional diversity loss alone yields no gain. Notably, combining IB with both the norm and diversity regularizers leads to a substantial improvement, reaching 49.1. This suggests that controlling the update scale (via trust-region) and promoting directional diversity must work together to fully exploit the potential of $\Delta$. This aligns with our design intuition: the trust-region constraint controls the update scale (via $\varepsilon$) but does not regulate directionality, which can lead to redundant or entangled corrections. The diversity regularization ensures that increments associated with the same anchor are directionally distinct, encouraging more uniform and structured behavior around each anchor.
@@
Context Modality Choice in $\psi$. We evaluate the effect of applying $\Delta_{ij}$ to either modality across datasets with contrasting characteristics. As shown in the left of Table 3, on MSR-VTT, injecting $\Delta$ on the text side ($t_{i}+\Delta_{ij}$) achieves the best performance, while applying it to the video side degrades results (R@1 drops to 47.4). This aligns with the fact that MSR-VTT contains many visually similar short clips, making it more effective to adjust the abstract text representation for fine-grained distinctions. Conversely, on ActivityNet, applying $\Delta$ to the video side leads to a notable performance boost, whereas modifying text harms results (R@1 $\approx$ 40). This is likely because the long but redundant videos are paired with rich, structured captions—making video-side adaptation more beneficial. These trends highlight the importance of aligning $\psi$’s modality choice with dataset structure. For DiDeMo, given its moderate video length (30s), we treat video as the anchor and set $c=v_{j}$ accordingly.
More ablation analysis including the lower bound $\lambda$ of $\mathcal{L}_{\varepsilon}$, the combination between context modality type $c$ of $\psi$ and $\eta$, the scale factor $\sigma$ of $\mathcal{L}_{\text{dir}}$ and the choice of anchor of $\mathcal{L}_{\text{IB}}$ can be found in Appendix.B.
5 Qualitative Analysis
To understand the effect of the learned increment $\Delta$, we conduct a qualitative analysis on the MSR-VTT 1k-A validation set, focusing on its impact on representation geometry and alignment behavior at inference. As shown in Figure 2(a), our method GARE yields lower cosine similarities between positive pairs than the baseline, indicating larger angular separation and improved uniformity on the unit hypersphere [40]. Figures 2(b) and 2(c) further show that while $t_{i}$ is concentrated around norm 12, the offset $t_{\Delta_{ij}}$ exhibits greater and more diverse magnitudes. This supports our view that $\Delta$ serves as a tension offloading mechanism, allowing $t_{i}$ to remain a stable prototype while $t_{\Delta_{ij}}$ absorbs pair-specific adjustments—particularly useful under gradient noise from false negatives. As shown in Figure 2(d), $t_{\Delta_{ij}}$ is generally farther from $v_{j}$ than $t_{i}$, indicating that the model does not simply pull pairs closer but expands the representation scale to allow more flexible alignment. This explains why, despite lower cosine similarity, GARE achieves better uniformity and geometric balance. Additional training-time analysis is provided in Appendix.C.
@@ 5 Qualitative Analysis
Figure 2: Qualitative analysis on the MSR-VTT 1k-A validation set. $t_{\text{new}}$ denotes $t_{\Delta}$. Our method induces greater angular separation between positive pairs (a), redistributes $t_{\Delta}$ norms to release gradient tension (b, c), and pushes $t_{\Delta}$ outward from $v_{j}$ (d), promoting uniformity.
@@ 6 Conclusion
This paper investigates the optimization behavior of contrastive learning in text-video retrieval, focusing on the tension caused by modality gap and the noise introduced by false negatives. We show that such tension arises when gradients from negative pairs conflict with positive updates, restricting the optimization freedom of text anchors. To address this, we perform a first-order Taylor expansion of the InfoNCE loss and derive a pair-specific increment $\Delta_{ij}$ as an optimal adjustment direction within a trust region. By learning $\Delta_{ij}$ through a neural module conditioned on the semantic gap, and regularizing its structure, we enable more stable and interpretable optimization. Experiments on standard retrieval benchmarks confirm the effectiveness of this approach, and qualitative results demonstrate that $\Delta$ captures meaningful geometry for resolving representation conflicts.
@@ Appendix A Limitations
A.1 Lack of Modality Alignment
The fundamental issue of modality gap persists. Despite reducing the tension between the embeddings of text and video, the two modalities still reside in completely disjoint regions of the representation space. Our method alleviates the modality gap by offloading some of the gradient tension, but it does not align the two modalities in a meaningful way. The solution we propose does not explicitly enforce a shared representation space between text and video; rather, it reduces the gradient tension through the optimization of $\Delta_{ij}$. As a result, the method mitigates the effects of the modality gap and reduces some noise, but it does not address the root cause of the misalignment between the two modalities.
A.2 Lack of Generalized Supervision for $\Delta$
@@
In summary, while the method presents an effective way of mitigating gradient tension and improving the performance of contrastive learning, these limitations highlight areas for further development in future work.
Appendix B More Ablation Experiments
@@
In the next four experiments conducted on the MSR-VTT-1KA dataset, we investigated the impact of different combinations of $c$ and $\eta$, the impact of the scale coefficient $\sigma$ in the directional diversity regularization, the impact of the lower bound of Norm-Based Regularization of Trust-Region Radii and the impact of the anchor choice of $\mathcal{L}_{\text{IB}}$.
B.1 Combination of Context Modality type $c$ and $\eta$ Choices
Table 4: Ablation on combinations of $c$ and $\eta$ (modality/context choice).
setting R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
$c=t_{i}$ $\eta=-1$ 46.6 74.2 83.8 2.0 12.9
$c=t_{i}$ $\eta=+1$ 48.2 74.2 83.2 2.0 12.9
$c=v_{j}$ $\eta=-1$ 47.8 74.1 82.8 2.0 12.4
$c=v_{j}$ $\eta=+1$ 49.1 74.7 83.6 2.0 12.0
As shown in Table 4, the results show: 1) Fine-tuning the text side is better: By comparing different configurations of $c$ and $\eta$, fine-tuning the text side consistently outperforms other choices. Specifically, when $c=t_{i}$ and $\eta=1$, the text representation adjusts the increment $\Delta$ effectively, allowing for a higher semantic distinction of the increment. 2) Choosing $v$ as the context for $\psi$ is better: In this experiment, choosing the video modality $v_{j}$ as the context for $\psi$ yielded better results, especially when $c=v_{j}$ and $\eta=1$, achieving the highest R@1 score. This suggests that learning more fine-grained representations from the video modality helps improve cross-modal interactions and boosts performance.
Since videos in the MSR-VTT dataset are often clips from the same source video, leading to a high number of similar video samples, fine-grained learning becomes crucial. Adding this fine-grained adjustment to the text side enables $\Delta$ to have better semantic distinction and enhances retrieval performance.
@@ B.1 Combination of Context Modality type $c$ and $\eta$ Choices
B.2 Ablation of the Regularization Scale Coefficient $\sigma$ in Directional Diversity
As shown in Table 5, the results indicate: 1) Smaller $\sigma$ (0.5): When $\sigma$ is small, the diversity of the delta vectors is limited, which weakens the model’s ability to learn fine-grained representations. Although the R@1 score is 47.6, the model’s ability to distinguish between samples is reduced due to limited directional diversity. 2) Moderate $\sigma$ (1.0 and 2.0): In these configurations, the directional diversity of delta vectors is more balanced, effectively enhancing the model’s ability to capture cross-modal variations.
Table 5: Ablation on scale coefficient $\sigma$ for directional diversity.
$\sigma$ R@1↑ R@5↑ R@10↑ MdR↓ MnR↓
0.5 47.6 74.5 83.4 2.0 12.1
1.0 48.5 74.4 83.9 2.0 12.1
2.0 49.1 74.7 83.6 2.0 12.0
3.0 46.2 74.4 82.9 2.0 12.1
4.0 47.0 74.3 83.4 2.0 12.1
5.0 47.2 74.1 82.8 2.0 12.4
The R@1 score improves to 49.1, showing that appropriate variance increases retrieval performance. 3) Larger $\sigma$ (3.0 and 5.0): When $\sigma$ is too large, the directional difference between delta vectors becomes excessive, introducing noise and decreasing model performance. Specifically, with $\sigma=3.0$, the R@1 score drops to 46.2.
From this experiment, it is evident that an appropriately moderate $\sigma$ enables delta vectors to have useful directional diversity, thereby improving cross-modal representation learning. However, excessively large $\sigma$ may result in noise that negatively impacts model performance.
B.3 Ablation on $\Delta$ Norm Regularization Strength

### v5

@@ 3.3 Ablative Analysis
Table 4: Ablation on Context Modality Choice of $\psi$. Text-to-video retrieval results on three datasets under different context modalities.
Dataset Context $\mathbf{C}$ R@1↑ R@5↑ R@10↑ MnR↓
MSR-VTT $\mathbf{T}_{\text{word}}$ 47.4 73.5 82.1 12.9
$\mathbf{V}_{\text{frame}}$ 49.1 73.3 82.2 12.4
ActivityNet $\mathbf{T}_{\text{word}}$ 42.6 73.6 84.4 6.8
$\mathbf{V}_{\text{frame}}$ 40.2 72.2 83.6 8.1
DiDeMo $\mathbf{T}_{\text{word}}$ 46.5 74.3 82.6 12.3
$\mathbf{V}_{\text{frame}}$ 47.6 75.4 83.1 12.0
Table 5: Ablation on the interaction mode of $\psi$ on Text-to-Video Retrieval results on MSR-VTT 1k-A. The variant removes the relative gap modeling by using $t_{i}$ as the query and $\mathbf{V}_{\text{frame}}$ as the key–value, producing $t_{ij}^{\prime}$ and $\Delta_{ij}=v_{j}-t_{ij}^{\prime}$. Our gap-aware design preserves pair-specific structure and yields superior alignment.
Interaction Mode of $\psi$ R@1↑ R@5↑ R@10↑ MnR↓
Query $t_{i}$ (no gap) 46.1 73.2 81.9 13.7
Query $v_{j}-t_{i}$ 49.1 74.7 83.6 12.0
Table 6: Ablation on the IB prior $r(\Delta)$ on MSR-VTT 1k-A. Comparison between normalized and unnormalized $\Delta_{ij}$ distributions with different Gaussian priors.
$\sigma$ R@1↑ R@5↑ R@10↑ MnR↓
Normalized $\Delta$
1.0 47.8 74.5 82.1 12.9
Unnormalized $\Delta$
0.1 47.7 73.4 82.2 12.9
1.0 49.1 74.7 83.6 12.0
10.0 48.1 74.6 83.5 12.0
100.0 48.6 74.7 83.2 11.8
Losses Combination.
@@
We conduct ablation studies on the MSR-VTT 1k-A validation set to assess the effectiveness of the proposed increment $\Delta$ and its associated regularizers. As shown in the right of Table 4, directly injecting $\Delta$ into the InfoNCE flow improves performance from 46.6 to 47.4, validating the benefit of gradient tension release via pairwise adjustment. Introducing the relaxed information bottleneck (IB) loss further boosts performance to 48.3, highlighting its role in guiding $\Delta$ toward semantically meaningful corrections. In contrast, adding the norm constraint or the directional diversity regularizer alone yields no gain, since each $\Delta_{ij}$ is only optimized with respect to its corresponding sample pair and thus lacks inherent contrastive behavior; simply minimizing InfoNCE can drive $\Delta$ toward trivial or collapsed solutions. The IB loss imposes a prior structure by restricting the optimization freedom of $\Delta$, and within this semantically grounded structure, the norm and diversity regularizers become truly effective; without such semantic constraints, applying them to trivial solutions of $\Delta$ would be meaningless. When combined, relaxed IB with norm and diversity regularization achieves the best performance (49.1), demonstrating that semantic grounding and structured regularization must work together to fully exploit the potential of $\Delta$.
Impact of Cross-attention Interaction Design.
As shown in Table 5, we examine the effect of replacing $\psi$’s pair-wise gap-aware interaction with a simplified query–key setting, where $t_{i}$ serves as the query and $\mathbf{V}{\text{frame}}$ as the key–value. This variant yields a fused representation $t{ij}^{\prime}$ and a residual $\Delta_{ij}=v_{j}-t_{ij}^{\prime}$, thereby removing explicit gap modeling between $t_{i}$ and $v_{j}$. Although it captures frame-level semantics aligned with $t_{i}$, directly updating $t_{i}$ via $\Delta_{ij}$ leads to severe performance degradation, as gradients near the loss side propagate to $t_{\Delta_{ij}}=t_{i}+v_{j}-t_{ij}^{\prime}$, disturbing the anchor semantics. To mitigate this, we normalize $\Delta_{ij}$ to a unit direction $\Delta_{\text{norm}}$ and compute a similarity-based magnitude $\mathcal{R}$ from frame-wise similarities between $t_{i}$ and $v_{j}$, obtained through a linear projection followed by exponentiation. The final correction $t_{\Delta_{ij}}=t_{i}+\mathcal{R}\cdot\Delta_{\text{norm}}$ partially alleviates gradient interference but still lacks explicit semantic-gap awareness, preventing $\psi$ from leveraging CLIP’s prior alignment—confirming that explicit pair-wise gap modeling is crucial for robust cross-modal alignment.
Prior $r(\Delta)$ of Information Bottleneck.
We further examine the effect of the prior distribution in the information bottleneck objective. A standard Gaussian prior is commonly used to regularize normalized embeddings, implicitly encouraging an isotropic distribution concentrated around a hyperspherical shell in high-dimensional space. In contrast, our best performance is achieved when the unnormalized $\Delta_{ij}$ distribution is regularized against the same standard Gaussian prior. As shown in Table 6, normalizing $\Delta_{ij}$ before KL regularization accelerates convergence but reduces final accuracy, as it constrains all increments to the unit sphere and removes one degree of freedom—preventing each $\Delta_{ij}$ from being regularized along its own radial axis. Moreover, since $\Delta_{ij}$ operates in the unnormalized space to correct the anchor $t_{i}$, preserving magnitude information is crucial for effective alignment. We further tested priors $r(\Delta)=\mathcal{N}(0,\sigma^{2}I)$ with different standard deviation $\sigma\in\{0.1,10,100\}$ under the unnormalized setting, but none outperformed the standard Gaussian ($\sigma=1$), suggesting that the optimal prior variance is data-dependent, while the standard Gaussian provides a balanced regularization strength in our case.
Context Modality Choice in $\psi$.
@@
We evaluate the effect of applying $\Delta_{ij}$ to either modality across datasets with contrasting characteristics. As shown in the left of Table 4, on MSR-VTT, injecting $\Delta$ on the text side (i.e., $t_{i}+\Delta_{ij}$) achieves the best performance, while applying it to the video side degrades results. This aligns with the fact that MSR-VTT contains many visually similar short clips, making it more effective to adjust the concise text representation for fine-grained distinctions. Conversely, on ActivityNet, applying $\Delta$ to the video side leads to a notable performance boost, whereas modifying text harms results (R@1 $\approx$ 40). This is likely because the long but redundant videos are paired with rich, structured captions—making video-side adaptation more beneficial. These trends highlight the importance of aligning $\psi$’s modality choice with dataset structure.
@@
Figure 3: Qualitative analysis on the MSR-VTT 1k-A validation set. $t_{\text{delta}}$ denotes $t_{\Delta}$. Our method induces greater angular separation between positive pairs (a), redistributes $t_{\Delta}$ norms to release gradient tension (b, c), and pushes $t_{\Delta}$ outward from $v_{j}$ (d), promoting uniformity.
3.4 Qualitative Analysis
Geometric Properties of $\Delta$.
To understand the effect of the learned increment $\Delta$, we conduct a qualitative analysis on the MSR-VTT 1k-A validation set, focusing on its impact on representation geometry and alignment behavior at inference. As shown in Figure 3(a), our method GARE yields lower cosine similarities between positive pairs than the baseline, indicating larger angular separation and improved uniformity on the unit hypersphere wang2020understanding .
Figures 3(b) and 3(c) show that the norm of the adjusted text embedding $t_{\Delta}$ consistently exceeds that of the original $t$, implying that $\Delta$ expands text representations onto a series of spheres of larger radius. Positive embeddings $t_{\Delta_{ii}}$ also have greater norms, consistent with our analysis of $\Delta$’s iterative update behavior. Performing the first-order Taylor expansion around nonzero $\Delta$ states (i.e., with nonzero initialization per iteration) mitigates logit-ranking distortion that occurs when expanding solely at zero. As negative increments $\Delta_{ij}$ ($j\neq i$) are pushed outward, the positive $\Delta_{ii}$ also expands to preserve relative belief masses sensoy2018evidential among logits, lowering overall cosine similarity while maintaining relative softmax probabilities.
Figure 3(d) further shows that $t_{\Delta_{ij}}$ lies farther from $v_{j}$ than $t_{i}$, implying that the model does not simply reduce inter-modal distance but expands the text representation onto a larger spherical shell for finer alignment. This suggests that, in cosine-based contrastive learning, encouraging greater dispersion of samples in the unnormalized feature space (i.e., higher pre-projection uniformity) may facilitate more effective alignment on the unit hypersphere.
Overall, these results indicate that optimization tension is effectively released: representation learning is no longer confined to the narrow region induced by the modality gap but occurs within a broader geometric space, thereby raising the upper bound of achievable alignment performance. Additional training-time analysis is provided in Appendix.G.
Gradient Analysis.
Refer to captionRefer to caption
Figure 4: Mean and variance of per-dimension gradients, indicating the positive gradients (top) acting on $t_{\Delta_{ii}}$ and $\Delta_{ii}$ and the sum of all negative gradients (bottom) for $t_{\Delta_{ij}}$ and $\Delta_{ij}$.
To further analyze how $\Delta$ redistributes optimization tension during training, we visualize the per-dimension gradient statistics of $t_{\Delta_{ij}}$ at a representative training step (Figure 4). In dimensions with significant optimization activity, both positive and negative gradients reach magnitudes around $g\approx 2.5$ and appear as approximate opposites, corresponding to the pair-specific gradient form in Eq. (8).
When aggregated over all pairs, these opposite gradients largely cancel out in the anchor’s update $\nabla_{t_{i}}\mathcal{L}_{i}$, leading to the near-zero gradient state described earlier. However, unlike $t_{i}$, which aggregates signals from all pairs, each $\Delta{ij}$ only receives the gradient transmitted from its corresponding pair. Consequently, the positive increment $\Delta_{ii}$ receives an effective gradient of approximately +g+g, while each negative $\Delta_{ij}$ receives around g/B-g/B, where $B$ denotes the batch size. Since these gradients act independently and are not mutually canceled, the total non-vanishing optimization strength per anchor $t_{i}$ is approximately $|\!+\!g|+B\!\cdot\!|\!-\!g/B|\approx 2|g|$, indicating that $\Delta$ components remain actively optimized rather than stagnant.
This reveals that the trajectory of $\Delta_{ij}$ reflects how $t_{i}$ explores the representation space. By distributing gradients across $\Delta$, our framework effectively offloads optimization tension from the anchor, expanding its reachable region and breaking the locality constraint imposed by the modality gap.
@@ 4 Conclusion
We revisited contrastive learning for text–video retrieval from an optimization perspective, identifying two key challenges: optimization tension from the modality gap and gradient noise from false negatives. Through a first-order Taylor expansion of the InfoNCE loss under a trust-region constraint, we derived a function space of increments $\Delta_{ij}$ that approximate descent directions. A learnable gap-aware network predicts $\Delta_{ij}$ to redistribute gradients across pairs, expanding the optimization range beyond the modality gap and mitigating false-negative noise. To ensure $\Delta_{ij}$ encodes compact yet informative corrections, we employ a variational information bottleneck with a relaxed compression objective for stable training. Experiments on four benchmarks demonstrate consistent improvements, and further analyses show that the learned increments form structured, semantically meaningful geometry that releases optimization tension and enhances alignment capacity.
5 Acknowlegements
This work was supported by the NSFC NO. 62172138 and No. 62202139. This work was also partially supported by the Fundamental Research Funds for the Central Universities NO. JZ2024HGTG0310 and No. JZ2025HGTB0226.
Appendix
This appendix provides additional discussions, derivations, and analyses that complement the main paper. We first discuss the limitations of our current framework and provide related works for broader context. Then, we present extended ablation studies and comparisons, followed by detailed mathematical derivations, including the relaxed form of our variational information bottleneck (VIB) objective and the rationale for using a non-zero state in the multivariate Taylor expansion. Finally, we provide implementation and efficiency details of the GARE model to support reproducibility and deployment.
@@ Appendix A Limitations
Lack of modality alignment.
The fundamental issue of modality gap persists. Despite reducing the tension between the embeddings of text and video, the two modalities still reside in completely disjoint regions of the representation space. Our approach alleviates this problem by releasing the optimization tension exerted on anchor representations, thereby expanding the effective optimization range of each text anchor through multiple pair-specific increments $\Delta_{ij}$. As a result, the method mitigates the effects of the modality gap and reduces some noise, but it does not address the root cause of the misalignment between the two modalities.
Lack of generalized supervision for $\Delta_{ij}$.
@@
Appendix B Related work
Contrastive learning and modality gap.
Contrastive learning has become a foundational paradigm in multimodal representation learning. Wang and Isola wang2020understanding formalize contrastive learning via the principles of alignment and uniformity on the hypersphere, offering a geometric perspective on representation quality. Wang and Liu wang2021understanding show that contrastive loss is hardness-aware and temperature-sensitive, and reveal a trade-off between representation uniformity and semantic tolerance, highlighting the need to preserve meaningful structure among semantically similar samples. Liang et al.liang2022mind investigate the modality gap in multimodal contrastive learning and attribute it to initialization imbalance and cone effects. False negatives have also been recognized as a key challenge in contrastive learning chuang2020debiased ; chen2021incremental , with solutions ranging from reweighting and elimination to dynamic detection and correction.
Text-video retrieval.
Text-video retrieval is one of the prominent tasks  Dong_2022 ; 9424429 ; Wang_2018_CVPR ; Baldrati_2022_CVPR ; he2023alignattendmultimodalsummarization ; xu2021videoclipcontrastivepretrainingzeroshot ; miech2020endtoendlearningvisualrepresentations ; 9879245 ; jin2023text ; Lee_2018_ECCV ; wang2023balanceactmitigatinghubness in cross-modal learning. The majority of existing research  li2023progressive ; fang2023uatvr ; jin2023diffusionret ; ma2022xclipendtoendmultigrainedcontrastive ; wang2023unifiedcoarsetofinealignmentvideotext ; bogolin2022crossmodalretrievalquerybank ; yang2024dgldynamicgloballocalprompt in this area utilizes a mapping technique that aligns both text and video inputs within a shared latent space to facilitate direct similarity assessment. CLIP4Clip luo2022clip4clip is the first to adapt CLIP radford2021learning for video-text retrieval via temporal frame aggregation. TS2-Net liu2022ts2 improves temporal modeling through token shift and selection. X-Pool gorti2022x introduces text-guided pooling to highlight salient video tokens. HBI jin2022expectation values possible correspondences between frames and words using Banzhaf interaction for sensitive and explainable cross-modal alignment. EMCL jin2022expectation introduces an expectation-maximization dempster1977maximum framework to learn a compact latent space where video and text features are represented as linear combinations of shared bases. This decomposition reduces the rank of the latent space, mitigating the modality gap and enhancing semantic alignment. Unlike prior works that refine matching structures, our method analyzes the gradient form of InfoNCE and introduces a learnable gap-aware increment $\Delta_{ij}$ to offload optimization tension, enabling structured optimization in a trust-region-aware formulation.
Appendix C More Ablation and Comparison Experiments
@@
In the next four experiments conducted on the MSR-VTT 1K-A dataset, we investigated the design of $\psi$ network inputs, the impact of the scale coefficient $\alpha$ and the orthogonality variant in the directional diversity regularization, the impact of the lower bound $\lambda$ of Norm-Based Regularization of Trust-Region Radii, and the impact of the anchor choice of $\mathcal{L}_{\text{IB}}^{\text{relax}}$. We also provide hard negative methods comparison.
C.1 Ablation on the Design of $\psi$ Network Inputs
Table 7: Ablation on context $\mathbf{C}$ and semantic gap $\eta$ under different correction modes. $\eta{=}1$ denotes vtv{-}t.
$\mathbf{C}$ $\eta$ R@1↑ R@5↑ R@10↑ MnR↓
t-corrected
$\mathbf{T}_{\text{word}}$ 1-1 46.6 74.2 83.8 12.9
$\mathbf{T}_{\text{word}}$ +1+1 48.2 74.2 83.2 12.9
$\mathbf{V}_{\text{frame}}$ 1-1 47.8 74.1 82.8 12.4
$\mathbf{V}_{\text{frame}}$ +1+1 49.1 74.7 83.6 12.0
v-corrected
$\mathbf{V}_{\text{frame}}$ 1-1 47.7 73.2 82.9 13.1
$\mathbf{V}_{\text{frame}}$ +1+1 47.5 73.5 81.8 13.0
$\mathbf{T}_{\text{word}}$ 1-1 48.2 73.6 82.8 12.7
$\mathbf{T}_{\text{word}}$ +1+1 48.8 73.2 83.6 12.9
We analyze the design of the $\psi$ network from three complementary aspects: 1) the context used for conditioning, 2) the semantic gap direction $\eta$, and 3) the corrected anchor (either $t$ or $v$). When correcting the text anchor ($t$), using the video-side sequence ($v$) as context yields better performance; conversely, when correcting the video anchor ($v$), using the text-side sequence ($t$) as context performs better. In both cases, the semantic gap is defined as vtv-t, suggesting that the query direction is asymmetric. Overall, performance is mainly determined by the interplay between the corrected anchor and the chosen context: cross-modal conditioning (e.g., correcting $t$ with $v$ as context) consistently outperforms uni-modal configurations, highlighting the benefit of cross-modal fusion.
Moreover, defining the semantic gap as vtv-t achieves better results than tvt-v. This can be attributed to the fact that video features tend to capture more shared and general concepts, while text features encode instance-specific information. Removing text-specific components from the video representation thus encourages more generalized and robust semantic alignment.
@@ C.1 Ablation on the Design of $\psi$ Network Inputs
C.2 Ablation of the Scale Coefficient $\alpha$ and the Orthogonality Variant in Directional Diversity
Table 8: Ablation on scale coefficient $\alpha$ for directional diversity.
$\alpha$ R@1↑ R@5↑ R@10↑ MnR↓
0.5 47.6 74.5 83.4 12.1
1.0 48.5 74.4 83.9 12.1
2.0 49.1 74.7 83.6 12.0
3.0 46.2 74.4 82.9 12.1
4.0 47.0 74.3 83.4 12.1
5.0 47.2 74.1 82.8 12.4
To further analyze the impact of the scale coefficient $\alpha$ in the directional diversity regularization, we conduct an ablation study as shown in Table 8. Recall that the loss adopts a log-mean-exp form:
$\mathcal{L}_{\text{dir}}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\Big[\log\mathbb{E}_{j,k}\left[\exp\left(-\alpha\cdot(1-\langle z_{ij},z_{ik}\rangle)\right)\right]\Big],$
where $z_{ij}=\frac{\Delta_{ij}}{|\Delta_{ij}|_{2}}$ denotes normalized directions under the same anchor $t_{i}$. The term $(1-\langle z_{ij},z_{ik}\rangle)$ lies in $[0,2]$, and the log-mean-exp operator favors pairs with higher cosine similarity (i.e., smaller angular distance), encouraging local angular diversity without enforcing strict orthogonality.
The coefficient $\alpha$ adjusts the strength of uniformity: 1) When $\alpha$ is small, the exponential term emphasizes pairs with higher similarity, focusing optimization on local clusters of $\Delta_{ij}$ directions. 2) As $\alpha$ increases, the optimization becomes more uniform across pairs, driving the distribution of $\Delta_{ij}$ directions toward isotropy.
Empirically, as shown in Table 8, performance peaks when $\alpha=2.0$, achieving the best R@1, R@5 and MnR. When $\alpha$ grows larger, R@1 and R@5 exhibit reduced variance, indicating that the model becomes less sensitive to angular differences. This aligns with our interpretation that large $\alpha$ values saturate the exponential term—causing $\exp(-\alpha(1-\langle z_{ij},z_{ik}\rangle))$ to approach zero uniformly—thus diminishing the discriminative effect among diverse directions. Overall, a moderate $\alpha$ achieves the best trade-off between directional diversity and training stability, effectively preventing directional collapse while maintaining meaningful variation across $\Delta_{ij}$.
To further enhance the distinction among $\Delta_{ij}$ directions, we additionally experimented with a strict orthogonalization objective:
$\mathcal{L}_{\text{orth}}=\frac{1}{B^{2}}\sum_{j}^{B}\sum_{k,k\neq j}^{B}\langle z_{ij},z_{ik}\rangle^{2},$
which enforces all direction pairs under each anchor $t_{i}$ to be mutually orthogonal. Unlike the log-mean-exp formulation, this loss treats both high- and low-similarity pairs equally and aims to drive all pairwise cosine similarities toward zero. However, this uniform orthogonality constraint yields inferior performance (R@1=47.7, R@5=74.3, R@10=83.1, MnR=12.1). We attribute this degradation to the excessive suppression of negatively correlated directions—pairs that are semantically opposite or far apart are also pushed toward zero similarity.